Some rends e you that hey paid $28.384 down on a new house and are o pay S811 per month or 30 years n e est s 6.6% comp our ed mor h what was he selling price of the house? How much interest wil he pay n 3 years Selling price of the house:s(Round to two decimal places as needed.)
Answer #1
1) Selling price of the house $ 1,55,368.86
Working:
# 1 Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.0055)^-360)/0.0055 i 6.6%/12 = 0.0055
= 156.5781 n 30*12 = 360
# 2 Present value of monthly cash flows = Cash flows * Present value of annuity of 1
= $                 811 * 156.5781
= $ 1,26,984.86
# 3 Down payment $     28,384.00
Present value of monthly cash flows $ 1,26,984.86
Present value of all cash flows $ 1,55,368.86
2) Interest paid $ 1,64,975.14
Working:
Total amount paid 811 * 30 * 12 = $ 2,91,960.00
Less:Present value of monthly payment $ 1,26,984.86
Interest paid $ 1,64,975.14